In pure copper at room temperature there is about one atom vacancy every 10^{10} atom sites.

At the metling temperture (1083 centigrade), there is about one vacanvy every 10^{3} sites.

Question:

What fraction of the volume change that takes place when the crystal is heated from room temperture to its metling temperature is due to the presence of vacancies?

The linear thermal expansion coefficient of copper is 16.5* 10^{-6} Centigrade^{-1}
Answer #1

Q1.

Fraction of Volume change = (Na*(1+x)/4)*Vc

Fraction of Volume change = (Na*(1+x)/4)*Vc

V1083 = (Na*(1+10^-3)/4)*Vc

solve

((V1083-Vrt)/Vrt=10^-3)

fraction --> 1.913*10^-2